OpenStudy (anonymous):

Evaluate the integral ((lnx)^3)/x dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?} ((\ln x)^{3})/x dx\]

OpenStudy (jtvatsim):

try a u-sub. :)

OpenStudy (anonymous):

I tried I got one result, it's..

OpenStudy (anonymous):

put lnx = t so (1/x)=dt

OpenStudy (anonymous):

then it will become simple

OpenStudy (anonymous):

\[\ln \left| x \right| \frac{ \ln x ^{4} }{ 4 } +C\]

OpenStudy (anonymous):

since the antiderivative of 1/x dx=ln abs(x) +C

OpenStudy (anonymous):

is this correct, or did I make a mistake?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

cool ^_^

OpenStudy (anonymous):

thanks!